OpenStudy (anonymous):

I do not understand how to do this equation x^2+14x+4=0

OpenStudy (anonymous):

It's an ugly one, but you can use the quadratic formula for solving quadratic equations, assuming you have learned this (and you should have).

OpenStudy (amistre64):

and and subtract 49 from it to complete the square

OpenStudy (anonymous):

Yep. As a reminder, for \(ax^2+bx+c = 0\), the roots are given by the quadratic formula, which you should memorize by heart.\[ x= \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Just one silly question that has nothing to do with the question, how do you get the right side expression to be a fraction like that? When I try, x comes in with it :P

OpenStudy (anonymous):

(that is how do you do that with the equation option)

OpenStudy (amistre64):

x^2 + 14x + 4 = 0 x^2 + 14x = -4 x^2 + 14x + 7^2 = -4 + 7^2 (x+7)^2 = 45 x+7 = +- sqrt(45) x = -7 +- sqrt(45)

OpenStudy (amistre64):

using the latex \frac{top}{bottom}

OpenStudy (amistre64):

\[\frac{top}{bottom}\]

OpenStudy (anonymous):

The derivation of the quadratic formula is rather sneaky. There are many places it's online -- I don't remember it 100%. It's too long to derive every time you want to get roots.

OpenStudy (anonymous):

Thanks amistre64, I have been longing for that!

OpenStudy (amistre64):

\[\text{\ [\frac{top}{bottom}\ ]}\] no spaces of course

OpenStudy (anonymous):

I see i have to divide by 2 with 14 to get 7

OpenStudy (amistre64):

yes, divide the 14 in half, then square it

OpenStudy (amistre64):

its called completeing the square

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you for helping me to understand this i need to practice more

OpenStudy (amistre64):

practice does help ;)

OpenStudy (amistre64):

completing the square is how you prove the quadratic formula

OpenStudy (anonymous):

the answer to the problem is -7-3sqrt 5and -7+3sqrt5, how did they get this answer

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[x^2+\frac{b}{a}x+\frac{c}{a}=0\] \[x^2+\frac{b}{a}x+(\frac{b}{2a})^2-(\frac{b}{2a})^2+\frac{c}{a}=0\] \[(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)-\frac{b^2}{4a}+\frac{4c}{4a}=0\] \[(x+\frac{b}{2a})^2=\frac{b^2}{4a^2}-\frac{4c}{4a^2}\] \[(x+\frac{b}{2a})^2=\frac{b^2-4ac}{4a^2}\] \[x+\frac{b}{2a}=\pm\sqrt{\frac{b^2-4ac}{4a^2}}\] \[x+\frac{b}{2a}=\pm\frac{\sqrt{b^2-4ac}}{2a}\] \[x=-\frac{b}{2a}\pm\frac{\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-14\pm\sqrt{14^2-4(1)(4)}}{2(1)}\] \[x=\frac{-14\pm\sqrt{196-16}}{2}\] \[x=\frac{-14\pm\sqrt{180}}{2}\] \[x=\frac{-14\pm\sqrt{9*4*5}}{2}\] \[x=\frac{-14\pm3*2\sqrt{5}}{2}\] \[x=\frac{-7\pm3\sqrt{5}}{1}\]

OpenStudy (anonymous):

ok I think i may understand now

OpenStudy (anonymous):

amistre64 - use \left( and \right) as parenthesis when you type your LaTeX to make them taller. ;-)

OpenStudy (anonymous):

ok